OpenStudy (anonymous):

-6x+2y=-6 , find slope intercept form. please help i dont get it

OpenStudy (abbles):

Ok Slope Intercept Form is y = mx + b

OpenStudy (abbles):

So you would need to isolate the y variable... do you know how to do that?

OpenStudy (abbles):

-6x+2y=-6 subtract the x and it's coefficient 2y = -6x - 6 divide by 2 y = -3x -2 And there you go, that is in Slope Intercept form.

OpenStudy (anonymous):

yes and i get -3

OpenStudy (anonymous):

Shouldn't it be +6x?

OpenStudy (anonymous):

-6x+2y-=-6 +6x on both sides 2y=6x-6 divide by 2 y=3x-3.

OpenStudy (abbles):

Oops yep that was a mistake. It would be positive 6 and so the answer would be: y = 3x - 3

OpenStudy (anonymous):

thank you so much abbles and halorazer

OpenStudy (abbles):

np :) thanks for the metal

OpenStudy (anonymous):

your welcome